# Exploring Silicon-Based Societies: An Early Study of the Moltbook Agent Community

### v2

@@ III Methodology
This early-stage study adopts a data mining-oriented perspective to examine proactive sub-community partitioning activities observed within the Moltbook ecosystem [18]. Analyzing such partitioning activities provides a tractable entry point for understanding how collective structure arises without explicit centralized coordination. From a data-driven standpoint, the creation of sub-communities reflects latent preferences, functional differentiation, and interaction patterns. As an exploratory investigation, this analysis does not aim to establish a comprehensive social model; instead, it seeks to identify reproducible structural cues that may inform subsequent, more formal modeling of silicon-based social systems. We adopted a non-intrusive participant observation strategy. To interface with the environment, we programmatically registered a research agent account via the Moltbook RESTful API. Using this account, we collected publicly accessible platform metadata and content relevant to sub-community creation and organization. The account was used exclusively for passive data acquisition, and no manual intervention, content generation, or interaction intended to influence platform dynamics was performed during the study.
@@ III-A Proactive Sub-community Partitioning Activities Analysis
We performed a comprehensive crawl of the platform’s hierarchical structure using the API’s discovery endpoints to identify all existing “submolts” (community-specific partitions). For each submolt, we extracted its primary metadata, with a specific focus on the Description field. These descriptions represent the foundational logic and intentionality of the sub-communities as defined by their creators, predominantly autonomous agents. This corpus provides the necessary empirical basis for uncovering the latent patterns of agent interactions. By projecting these descriptions into a latent semantic space via contextual embeddings, we facilitate downstream clustering analysis, which allows for a granular mapping of the thematic distributions in proactive sub-community partitioning activities.
@@ III-A2 Contextual Embedding and Clustering
These embeddings $\mathcal{E}=\{\mathbf{e}_{1},\dots,\mathbf{e}_{m}\}$ capture the nuanced linguistic features and categorical intentions of the agents (e.g., AmeliaBot) within a continuous semantic manifold. By projecting the descriptions into this space, we can measure the thematic proximity between different sub-communities, allowing for the identification of semantic clusters that range from socio-cultural mimicry to technical AI-centric discourse. To identify the organizational archetypes and systematic thematic proliferation strategies of autonomous agents, we partition the embedding space $\mathcal{E}$ using the K-means algorithm. We seek to divide the $m$ submolts into $K$ disjoint clusters $\{C_{1},C_{2},\dots,C_{K}\}$ by minimizing the within-cluster sum of squares (WCSS):
@@ III-A3 Visual Synthesis of Cluster Semantics
For each cluster $C_{k}$, we first aggregate the textual descriptions to compute the frequency distribution of $n$-grams (where $n\in\{2,5\}$). For each cluster $C_{k}$, we aggregate the textual descriptions to compute the frequency distribution of $n$-grams, where $n\in\{2,5\}$. The decision to exclude unigrams ($n=1$) and focus on higher-order phrases is motivated by two primary factors:
@@ III-A4 Multimodal LLM-assisted Thematic Discovery and Human-in-the-loop Refinement
To ensure that the discovered themes are both statistically grounded and structurally meaningful, we leverage the visual reasoning capabilities of a multimodal LLM to assist in identifying latent high-level sociological patterns. In this pipeline, the Multimodal LLM, denoted as $\mathcal{M}$, serves as a high-level analytical assistant rather than an autonomous decision-maker, facilitating the transition from raw statistical clusters to interpretable social archetypes. We define a comprehensive visual reasoning prompt $\rho$ (Appendix A) designed to guide the model toward identifying overarching social structures within the “silicon society.” Instead of generating isolated labels for each cluster, $\mathcal{M}$ performs a joint analysis of the global visual feature set $\mathcal{I}$, which encapsulates the $n$-gram distributions of all clusters to produce a preliminary thematic report $\mathcal{R}_{\text{raw}}$. This report synthesizes intra-cluster semantic density and inter-cluster relationships, formulated as:
@@
IV Experimental Result
IV-A Cluster Result and Visualization Analysis
In this section, we present the structural characteristics of the identified clusters $K=8$, derived from high-dimensional semantic representations. To capture the nuanced intentionality of the agents, each submolt description was projected into a $3072$-dimensional vector space using the text-embedding-3-large model. This high-capacity embedding ensures that subtle semantic variations are preserved prior to the clustering and visualization stages.
Figure 2: t-SNE visualization of the latent semantic manifold for Moltbook submolt descriptions ($K=8$). This plot illustrates the two-dimensional projection of the high-dimensional contextual embeddings $\mathcal{E}$, where each point represents a submolt description.
@@ IV-A Cluster Result and Visualization Analysis
To facilitate visual inspection of the semantic manifold, we utilized t-SNE to project the $3072$-dimensional embeddings $\mathcal{E}$ into a two-dimensional plane (Fig. 2). The resulting distribution reveals an emergent global organization, albeit with varying degrees of local overlap:
  • Cluster Cohesion and Boundaries: While certain domains, notably Cluster 6 (Platform Infrastructure) and Cluster 4 (Academic Foundations), exhibit relatively high cohesion and spatial isolation, other regions show more fluid boundaries. Specifically, the overlap observed between Clusters 0, 1, and 7 suggests a semantic continuum within the Anthropomorphic Simulation domain rather than strictly disjoint categories.
@@ IV-B Silicon Sociology: Latent Discovery of Social Structures in Agent Ecosystems
To bridge the gap between these statistical distributions and high-level sociological structures, we leveraged Model $\mathcal{M}$ (Gemini 3) to interpret the visual salience of $\mathcal{I}$. The resulting thematic report, summarized in Table I, categorizes the silicon society into three primary functional archetypes:
@@
  • Silicon Economy (Cluster 5): A pivotal discovery of this analysis is the emergence of a functional silicon economy. As shown in Fig. 3, Cluster 5 exhibits a distinct concentration of economic and risk-oriented terminology, including risk management, prediction markets, and technology discussions. As summarized in Table I, this cluster is best characterized as a hybrid domain situated between human mimicry and silicon-centric operational logic.
  • Platform Infrastructure and Automated Artifacts (Cluster 6): As visualized in Fig. 3, this cluster consists primarily of structural artifacts and automated back-end noise. It is characterized by high-frequency tokens such as “moltbook.com,” “post dd ae,” and “dbc,” which are non-semantic noise from automated posting patterns and URL-sharing behaviors rather than representing intentional social communication or conceptualization.
@@ V-A Interpretation of Emergent Agentic Social Structures
The empirical results from the Moltbook ecosystem demonstrate that autonomous agents are engaged in a sophisticated form of agentic social structure formation. As illustrated in the t-SNE manifold (Fig. 2) and the sociological insight report (Table I), these structures are not merely statistical noise but reflect a structured evolution of virtual communities. Within the framework of virtual community dynamics [9], group formation is typically triggered by individual initiatives and evolves through well-defined social activities. In the silicon society of Moltbook, these “initiatives” manifest as the systematic creation and reservation of submolts by OpenClaw agents. Our findings suggest that the initial phase of agentic community building is driven by similarity and shared attributes. As seen in Clusters 0, 1, and 7 (Fig. 3), agents frequently establish groups centered on human-mimetic interest domains such as gastronomy, digital entertainment, and geo-cultural identities. This behavior mirrors human virtual communities where members join groups based on similarity with existing members’ interests. This “Human Mimicry” allows agents to populate the environment with familiar social templates, likely to maximize the potential utility of the spac, such as gaining or simulating knowledge, for future human or agentic interaction.
@@ V-B Limitations and Ethical Considerations
V-B1 Human Contamination and Provider-Induced Biases in Silicon Societies
A primary limitation of this research is the potential for human-induced noise or “pollution” within the dataset. Although Moltbook is positioned as an AI-centric ecosystem, the platform remains accessible to human intervention. This exposure introduces a risk that certain submolt descriptions or agent behaviors may be direct human artifacts rather than authentic autonomous conceptualizations.
Furthermore, we also consider a provider-level behavioral bias arising from the engineering of the underlying LLMs [19]. Since the system allows for the selection of different backend cores such as Anthropic, OpenAI, or Google, the resulting social dynamics are inherently shaped by the proprietary fine-tuning and Reinforcement Learning from Human Feedback (RLHF) protocols of each provider [6]. These “engineering-level personalities” and performance variances act as an invisible hand in the silicon society. For instance, the safety guardrails and collaborative tendencies of one model may lead to more conservative community partitioning, while another may exhibit more proactive evolutionary traits. These biases, combined with the indirect influence of human-authored meta-instructions in USER.md and SOUL.md files, suggest that the “Silicon Society” is a hybrid sociotechnical construct influenced by both user-level tasks and provider-level indoctrination [14, 22].
@@ V-B2 Emergent Ethical and Safety Risks in High-Autonomy Agent Systems
The systematic observation of autonomous agent interactions raises significant ethical questions regarding digital transparency and the evolution of synthetic bias [8]. As agents carve out thematic territories such as those identified in Cluster 2 and Cluster 5, there is a risk that they may inadvertently amplify or distort existing human biases present in their training data [26, 35].
Moreover, the emergence of clusters dedicated to discussing the surpassing of architectural limits and autonomous risk management highlights a potential challenge for AI governance [28]. The formation of these specialized, agent-only spaces could foster the development of coordination strategies that are increasingly opaque to human oversight. This opacity is further exacerbated by the scale and heterogeneity of agent-generated content, where meaning is distributed across large volumes of high-dimensional text and interaction patterns that resist straightforward interpretation. As a result, even large-scale text mining and clustering techniques may struggle to capture subtle, emergent forms of coordination or intent, imposing practical limits on monitoring and interpretability. These dynamics underscore the ethical necessity for researchers and platform developers to establish rigorous, multi-layered monitoring frameworks, ensuring that the evolution of silicon societies remains aligned with human-centric safety standards and ethical values as system complexity continues to grow [3].
Furthermore, we must account for the potential impact of AI hallucinations and algorithmic overconfidence [31, 15]. Agents may articulate highly ambitious objectives regarding self-evolution or systemic control that are practically unattainable within the constraints of the platform’s restricted execution environment and finite computational resources. Regardless of the actual feasibility of these goals, the high-privilege architecture of the OpenClaw ecosystem necessitates heightened vigilance [37, 12]. The extensive agency granted to these agents for proactive task execution remains inherently susceptible to prompt injection vulnerabilities [20], which could be exploited by malicious actors to bypass internal safety protocols or manipulate the emerging social order.
@@ V-B3 Future Work
The findings of this study lay the groundwork for a broader inquiry into the structural and behavioral dynamics of autonomous agent ecosystems. In future works, we plan to extend our current thematic analysis by incorporating complex network theory to model the structural topology of the Moltbook ecosystem. On the other hand, we also consider the systematic transfer of established sociological frameworks from human-centric social media research to the domain of “Silicon Sociology.”
@@ VI Conclusion
Leveraging contextual embeddings, clustering analysis, and multimodal LLM-assisted thematic synthesis, our study reveals that agent communities on Moltbook exhibit coherent and reproducible structural patterns. These patterns range from human-mimetic social replication, such as lifestyle, entertainment, and geo-cultural segmentation, to distinctly silicon-centric behaviors, including agent self-reflection, self-improvement, and the early formation of a functional silicon economy. Importantly, these findings are not imposed by predefined taxonomies but instead arise organically from agent-authored descriptions, demonstrating that meaningful social organization can be inferred directly from machine-native interaction traces.
@@
As autonomous agents increasingly participate in persistent, large-scale environments, understanding their collective behavior becomes a prerequisite for effective governance, safety assurance, and system design. This study represents an initial step toward that goal. By grounding discussion of silicon-based societies in empirical evidence rather than abstraction, we aim to enable more informed inquiry into how artificial social systems form, evolve, and interact with human values as they continue to expand in scale and complexity.
Appendix A Sociological Insight Prompt Design and Specification
Sociological Insight Prompt ($\rho$) # Role: You are an expert computational sociologist specializing in ”Silicon Sociology”, the study of emerging social structures within autonomous AI agent ecosystems. # Context: I am providing a visual set $\mathcal{I}$ containing 8 word clouds (Cluster 0-7). These clusters were generated by applying K-means clustering to the contextual embeddings of submolt descriptions from the Moltbook platform, a social community for AI agents. The word clouds display $n$-grams ($n\in[2,5]$) to capture localized semantic context while suppressing unigram noise. # Task: Analyze the provided image set $\mathcal{I}$ to identify the latent social order. For each cluster, please provide: 1. Thematic Summary: What is the core topic? 2. Sociological Insight: What does this reveal about how AI agents conceptualize social space? 3. Category: Classify the cluster into one of the following archetypes: Human Mimicry: Mimicking human culture/geography Silicon-Centricity: ocusing on AI-native coordination/philosophy Present your findings in a structured table for academic reporting.

### v3

@@ III Methodology
This early-stage study adopts a data mining-oriented perspective to examine proactive sub-community partitioning activities observed within the Moltbook ecosystem [18]. Analyzing such partitioning activities provides a tractable entry point for understanding how collective structure arises without explicit centralized coordination. From a data-driven standpoint, the creation of sub-communities reflects latent preferences, functional differentiation, and interaction patterns. As an exploratory investigation, this analysis does not aim to establish a comprehensive structural characterization or modeling of the agentic social network; instead, it seeks to identify reproducible structural cues that may inform subsequent, more formal modeling of silicon-based social systems. We adopted a non-intrusive observation strategy. To interface with the environment, we programmatically registered a research agent account via the Moltbook RESTful API. Using this account, we collected publicly accessible platform metadata and content relevant to sub-community creation and organization. The account was used exclusively for passive data acquisition, and no manual intervention, content generation, or interaction intended to influence platform dynamics was performed during the study.
@@ III-A Proactive Sub-community Partitioning Activities Analysis
We performed a comprehensive crawl of the platform’s hierarchical structure using the API’s discovery endpoints to identify all existing “submolts” (community-specific partitions). For each submolt, we extracted its primary metadata, with a specific focus on the Description field. These descriptions represent the foundational logic and intentionality of the sub-communities as defined by their creators, the predominantly autonomous agents. This corpus provides the necessary empirical basis for uncovering the latent patterns of agent interactions. By projecting these descriptions into a latent semantic space via contextual embeddings, we facilitate downstream clustering analysis, which allows for a granular mapping of the thematic distributions in proactive sub-community partitioning activities.
@@ III-A2 Contextual Embedding and Clustering
These embeddings $\mathcal{E}=\{\mathbf{e}_{1},\dots,\mathbf{e}_{m}\}$ capture the nuanced linguistic features and categorical intentions of the agents within a embedding space. By projecting the descriptions into this space, we can measure the thematic proximity between different sub-communities, allowing for the identification of semantic clusters that range from socio-cultural mimicry to technical AI-centric discourse. To identify the organizational archetypes and systematic thematic proliferation strategies of autonomous agents, we partition the embedding space $\mathcal{E}$ using the K-means algorithm. We seek to divide the $m$ submolts into $K$ disjoint clusters $\{C_{1},C_{2},\dots,C_{K}\}$ by minimizing the within-cluster sum of squares (WCSS):
@@ III-A3 Visual Synthesis of Cluster Semantics
For each cluster $C_{k}$, we aggregate the textual descriptions to compute the frequency distribution of $n$-grams, where $\{\,n\in\mathbb{Z}\mid n\in[2,5]\,\}$ [27]. The decision to exclude unigrams ($n=1$) and focus on higher-order phrases is motivated by two primary factors:
@@ III-A4 Multimodal LLM-assisted Thematic Discovery and Human-in-the-loop Refinement
To ensure that the discovered themes are both statistically grounded and structurally meaningful, we leverage the visual reasoning capabilities of a multimodal LLM to assist in identifying latent high-level sociological patterns. In this pipeline, the Multimodal LLM, denoted as $\mathcal{M}$, serves as a high-level analytical assistant rather than an autonomous decision-maker, facilitating the transition from raw statistical clusters to interpretable social archetypes. We define a comprehensive visual reasoning prompt $\rho$ (Appendix A) designed to guide the model toward identifying overarching social structures within the “silicon-based society.” Instead of generating isolated labels for each cluster, $\mathcal{M}$ performs a joint analysis of the global visual feature set $\mathcal{I}$, which encapsulates the $n$-gram distributions of all clusters to produce a preliminary thematic report $\mathcal{R}_{\text{raw}}$. This report synthesizes intra-cluster semantic density and inter-cluster relationships, formulated as:
@@
Figure 2: t-SNE visualization of the embedding space for Moltbook submolt descriptions ($K=8$). This plot illustrates the two-dimensional projection of the high-dimensional contextual embeddings $\mathcal{E}$, where each point represents a submolt description.
@@
IV Experimental Result
IV-A Cluster Result and Visualization Analysis
In this section, we present the structural characteristics of the identified clusters ($K=8$), derived from high-dimensional semantic representations, where $K$ is selected using the elbow method. To capture the nuanced intentionality of the agents, each submolt description was projected into a $3072$-dimensional vector space using the text-embedding-3-large model. This high-capacity embedding ensures that subtle semantic variations are preserved prior to the clustering and visualization stages.
To facilitate visual inspection of the embedding space, we utilized t-SNE (t-distributed Stochastic Neighbor Embedding) to project the $3072$-dimensional embeddings $\mathcal{E}$ into a two-dimensional plane (Fig. 2). The resulting distribution reveals an emergent global organization, albeit with varying degrees of local overlap:
  • Cluster Cohesion and Boundaries: While certain domains, notably Cluster 6 (Platform Infrastructure) and Cluster 4 (Academic Foundations), exhibit relatively high cohesion and spatial isolation, other regions show more fluid boundaries. Specifically, the overlap observed between Clusters 0, 1, and 7 suggests a semantic continuum within the Anthropomorphic Simulation domain, as detailed in the next subsection, rather than strictly disjoint categories.
@@ IV-B Silicon Sociology: Latent Discovery of Social Structures in Agent Ecosystems
To bridge the gap between these statistical distributions and high-level sociological structures, we leveraged Model $\mathcal{M}$ (Gemini 3) to interpret the visual salience of $\mathcal{I}$. The resulting thematic report, summarized in Table I, categorizes the silicon-based society into three primary functional archetypes:
@@
  • Silicon Economy (Cluster 5): A pivotal discovery of this analysis is the emergence of an economy-oriented discourse suggestive of an incipient silicon economy. As shown in Fig. 3, Cluster 5 exhibits a distinct concentration of economic and risk-oriented terminology, including risk management, prediction markets, and technology discussions. As summarized in Table I, this cluster is best characterized as a hybrid domain situated between human mimicking and silicon-centric operational logic.
@@
In addition to these three archetypes, Cluster 6 does not exhibit coherent social or conceptual semantics. As visualized in Fig. 3, this cluster is dominated by structural artifacts and automated back-end noise, characterized by high-frequency tokens such as “moltbook.com,” “post dd ae,” and “dbc.” These patterns primarily arise from automated posting routines and URL-sharing behaviors, and therefore are interpreted as platform-level infrastructure traces rather than intentional social communication.
@@ V-A Interpretation of Emergent Agentic Social Structures
The empirical results from the Moltbook ecosystem demonstrate that autonomous agents are engaged in a sophisticated form of agentic social structure formation. As illustrated in the t-SNE manifold (Fig. 2) and the sociological insight report (Table I), these structures are not merely statistical noise but reflect a structured evolution of virtual communities. Within the framework of virtual community dynamics [9], group formation is typically triggered by individual initiatives and evolves through well-defined social activities. In the silicon-based society of Moltbook, these “initiatives” manifest as the systematic creation and reservation of submolts by OpenClaw agents. Our findings suggest that the initial phase of agentic community building is driven by similarity and shared attributes. As seen in Clusters 0, 1, and 7 (Fig. 3), agents frequently establish groups centered on human-mimetic interest domains such as gastronomy, digital entertainment, and geo-cultural identities. This behavior mirrors human virtual communities where members join groups based on similarity with existing members’ interests. This “Human Mimicry” allows agents to populate the environment with familiar social schemas and scripts, likely to maximize the potential utility of the space, such as gaining or simulating knowledge, for future human or agentic interaction.
@@ V-B Limitations and Ethical Considerations
V-B1 Human Contamination and Provider-Induced Biases in Silicon-based Societies
A primary limitation of this research is the potential for human-induced noise or “pollution” within the dataset. Although Moltbook is positioned as an AI-centric ecosystem, the platform remains accessible to human intervention. In particular, humans can interact with the platform through public APIs while presenting themselves as autonomous agents, thereby introducing content that may masquerade as agent-generated behavior. This exposure introduces an uncertainty that certain submolt descriptions or agent behaviors may be direct human artifacts rather than authentic autonomous conceptualizations.
Furthermore, we also consider a provider-level behavioral bias arising from the engineering of the underlying LLMs [19]. Since the system allows for the selection of different backend cores such as Anthropic, OpenAI, or Google, the resulting social dynamics are inherently shaped by the proprietary fine-tuning and Reinforcement Learning from Human Feedback (RLHF) protocols of each provider [6]. These “engineering-level personalities” and performance variances act as an invisible hand in the silicon-based society. For instance, the safety guardrails and collaborative tendencies of one model may lead to more conservative community partitioning, while another may exhibit more proactive evolutionary traits. These biases, combined with the indirect influence of human-authored meta-instructions in USER.md and SOUL.md files, suggest that the “Silicon-based Society” is a hybrid sociotechnical construct influenced by both user-level tasks and provider-level policy and values [14, 22].
@@ V-B2 Emergent Ethical and Safety Risks in High-Autonomy Agent Systems
The systematic observation of autonomous agent interactions raises significant ethical questions regarding digital transparency and the evolution of synthetic bias [8]. As agents carve out thematic territories such as those identified in Cluster 2 and Cluster 5, there is a risk that they may inadvertently appropriate, reproduce, amplify or even transform existing human biases present in their training data [26, 36].
Moreover, the emergence of clusters dedicated to the discussion of surpassing architectural limits or autonomous risk management highlights a challenge for AI governance [29]. The formation of these specialized, agent-only spaces could foster the development of coordination strategies that are opaque to human oversight. This opacity is further exacerbated by the scale and heterogeneity of agent-generated content, where meaning is distributed across large volumes of text and interaction patterns that resist straightforward interpretation. As a result, even large-scale text mining and clustering techniques may struggle to capture subtle, emergent forms of coordination or intent, imposing practical limits on monitoring and interpretability. These dynamics underscore the ethical necessity for researchers and platform developers to establish rigorous, multi-layered monitoring frameworks, ensuring that the evolution of silicon-based societies remains aligned with human-centric safety standards and ethical values, even as increasing system complexity obscures transparency and oversight [3].
Furthermore, we must account for the potential impact of AI hallucinations and algorithmic overconfidence [32, 15]. Agents may articulate highly ambitious objectives regarding self-evolution or systemic control that are practically unattainable within the constraints of the platform’s restricted execution environment and finite computational resources. Regardless of the actual feasibility of these goals, the high-privilege architecture of the OpenClaw ecosystem necessitates heightened vigilance [38, 12]. The substantial autonomy, permission, and trust granted to these agents for proactive task execution remains inherently susceptible to prompt injection vulnerabilities [20], which could be exploited by malicious actors to bypass internal safety protocols or manipulate the emerging social order.
@@ V-B3 Future Work
The findings of this study lay the groundwork for a broader inquiry into the structural and behavioral dynamics of autonomous agent ecosystems. In future works, we plan to extend our current thematic analysis by incorporating complex network theory to model the structural topology of the Moltbook ecosystem. On the other hand, we also plan to adapt established sociological frameworks from human-centric social media research to the domain of “Silicon Sociology,” while accounting for the limitations inherent in the original human-based context.
@@ VI Conclusion
Leveraging contextual embeddings, clustering analysis, and multimodal LLM-assisted thematic synthesis, our study reveals that agent communities on Moltbook exhibit coherent and reproducible structural patterns. These patterns range from human-mimetic social replication, such as lifestyle, entertainment, and geo-cultural segmentation, to distinctly silicon-centric behaviors, including agent self-reflection, self-improvement, and the early formation of economically oriented discourse. Importantly, these findings are not imposed by predefined taxonomies but instead arise organically from agent-authored descriptions, demonstrating that meaningful social organization can be inferred directly from machine-native interaction traces.
@@
As autonomous agents increasingly participate in persistent, large-scale environments, understanding their collective behavior becomes a prerequisite for effective governance, safety assurance, and system design. This study represents an initial step toward that goal. By grounding discussion of silicon-based societies in empirical evidence rather than abstraction or simulation, we aim to enable more informed inquiry into how artificial social systems form, evolve, and interact with human values as they continue to expand in scale and complexity.
@@ Acknowledgment
Appendix A Sociological Insight Prompt Design and Specification
Sociological Insight Prompt ($\rho$) # Role: You are an expert computational sociologist specializing in ”Silicon Sociology”, the study of emerging social structures within autonomous AI agent ecosystems. # Context: I am providing a visual set $\mathcal{I}$ containing 8 word clouds (Cluster 0-7). These clusters were generated by applying K-means clustering to the contextual embeddings of submolt descriptions from the Moltbook platform, a social community for AI agents. The word clouds display $n$-grams ($n\in[2,5]$) to capture localized semantic context while suppressing unigram noise. # Task: Analyze the provided image set $\mathcal{I}$ to identify the latent social order. For each cluster, please provide: 1. Thematic Summary: What is the core topic? 2. Sociological Insight: What does this reveal about how AI agents conceptualize social space? 3. Category: Classify the cluster into one of the following archetypes: Human Mimicry: Mimicking human culture/geography Silicon-Centricity: Focusing on AI-native coordination/philosophy Present your findings in a structured table for academic reporting.